# Trees of Graphs as Boundaries of Hyperbolic Groups

### v1

@@ 2.3. Limit of a tree system
We now describe, following the exposition in Section 1 of [23], an object called the limit of a tree system $\Theta=(T,\{K_{t}\},\{\Sigma_{u}\},\{\varphi_{e}\})$, denoted $\lim\Theta$, starting with its description as a set. Denote by #$\#\Theta$ the quotient
@@
Again, for any finite b-subtree $F$ of $T$ let $R_{F}$ be the set of all rays $\rho$ in $T$ with $\rho(0)\in N_{F}$, and disjoint with $F$. Note that the map $\rho\to[\rho]$ is then a bijection from $R_{F}$ to the set $\partial T$ of all ends of $T$. Given a subset $U\subset K_{F}$ as above (i.e. saturated with respect to $\Omega_{F}$), put
@@
where all summands in the above union are viewed as subsets of $\lim\Theta$; clearly, ${\mathcal{G}}(U)$ is then also a subset in $\lim\Theta$. We consider the topology in the set $\lim\Theta$ given by the basis $\mathcal{B}$ consisting of all sets ${\mathcal{G}}(U)$, for all finite b-subtrees $F$ in $T$, and all open subsets $U\subset K_{F}$ saturated with respect to $\Omega_{F}$. We leave it as an easy and instructive exercise to check that the family $\mathcal{B}$ is closed under finite intersections, and hence it satisfies the axioms for a basis of topology. Moreover, we have the following result, which is a slight extension of Proposition 1.C.1 of [23], and which can be justified by the same proof as the latter proposition.
@@ 2.4. Functoriality with respect to subtrees
Next observation follows by a direct argument (compare the proof of Claim 1 inside the proof of Theorem 3.A.1 in [23]).
@@
In the case when $S=F$ is a finite b-subtree of $T$, it is not hard to see that the limit $\lim\Theta_{F}$ coincides with the partial union $K_{F}$. It follows then from Fact 2.5 that each of the partial unions $K_{F}$ naturally embeds (as a topological space) in $\lim\Theta$. In particular, each of the constituent spaces $K_{t}$ naturally embeds in $\lim\Theta$. A similar observation shows that each of the peripheral spaces $\Sigma_{u}$ naturally embeds in $\lim\Theta$. The following result is proved as Proposition 1.C.1 in [23].
@@ 2.5. An isomorphism of tree systems
Let $\Theta=(T,\{K_{t}\},\{\Sigma_{u}\},\{\varphi_{e}\})$ and $\Theta^{\prime}=(T^{\prime},\{K^{\prime}_{t}\},\{\Sigma^{\prime}_{u}\},\{\varphi^{\prime}_{e}\})$ be two tree systems of metric compacta. An isomorphism $F:\Theta\to\Theta^{\prime}$ is a tuple $F=(\lambda,\{f_{t}\},\{g_{u}\})$ such that:
@@
If $\Theta,\Theta^{\prime}$ are isomorphic tree systems of metric compacta then their limits $\lim\Theta$ and $\lim\Theta^{\prime}$ are homeomorphic.
@@ 2.6. An inverse system associated to a tree system
be a tree system. For any finite b-subtree $F\subset T$, consider the family $\Omega_{F}$ of peripheral subspaces of the partial union $K_{F}$, as described in Subsection 2.3, and denote by $K_{F}^{*}=K_{F}/\Omega_{F}$ the quotient of $K_{F}$ in which all the sets from $\Omega_{F}$ are shrinked to points. More precisely, let ${\mathcal{D}}_{F}$ be the decomposition of $K_{F}$ consisting of the sets of $\Omega_{F}$ and the singletons from the complement $K_{F}\setminus\bigcup\Omega_{F}$, and let $K_{F}^{*}=K_{F}/{\mathcal{D}}_{F}$. Since the family $\Omega_{F}$ is null, the decomposition ${\mathcal{D}}_{F}$ is upper semicontinuous, and hence the quotient $K_{F}^{*}$ is metrizable (see [9], Proposition 3 on page 14 and Proposition 2 on page 13). Consequently, $K_{F}^{*}$ is a metric compactum, and we call it the reduced partial union of $\Theta$ related to $F$. We denote by $q_{F}:K_{F}\to K_{F}^{*}$ the quotient map resulting from the above description of $K_{F}^{*}$.
@@
Next result is a slight extension of Proposition 1.D.1 in [23], and its proof is the same as the proof of that proposition (hence we omit it).
@@ 2.7. Consolidation of a tree system
Let $\Theta=(T,\{K_{t}\},\{\Sigma_{u}\},\{\varphi_{e}\})$ be a tree system of metric compacta. Let $\Pi$ be a partition of a tree $T$ into b-subtrees, i.e. a family of b-subtrees $S\subset T$ such that the black vertex sets $V_{S}^{b}:S\in{\Pi}$ are pairwise disjoint and cover all of $V_{T}^{b}$. We allow that some of the subtrees $S\in{\Pi}$ are just single black vertices of $T$.
@@
described as follows. As a set $V^{b}_{T^{\Pi}}$ of black vertices of $T^{{\Pi}}$ we take the family $\Pi$, and as a set $V^{w}_{T^{\Pi}}$ of white vertices we take the set of those white vertices $u\in V^{w}_{T}$ which remain outside all of the subtrees $S\in\Pi$. As the edge set $E_{T^{\Pi}}$ we take the set $\{e\in E_{T}:e\notin\cup_{S\in{\Pi}}\,E_{S}\}$, so that each edge $e$ from this set is meant to connect the white vertex $w(e)\in V^{b}_{T^{\Pi}}$ with this subtree $S\in\Pi$ (viewed as a black vertex of $T^{\Pi}$) which contains the vertex $b(e)$ of $T$. Intuitively speaking, the above described tree $T^{\Pi}$ is a “dual tree” of the partition $\Pi$.
To desribe the constituent spaces $K_{S}$ of the tree system $\Theta^{\Pi}$, for any subtree $S\in{\Pi}$ denote by $\Theta_{S}$ the restriction of $\Theta$ to $S$, and put $K_{S}:=\lim\Theta_{S}$. To describe the peripheral spaces $\Sigma_{u}^{\Pi}$ and the connecting maps $\varphi_{e}^{\Pi}$, note that for any vertex $u\in V^{w}_{T^{\Pi}}$ (viewed as a white vertex of $T$) and any edge $e$ with $w(e)=u$, the map $\varphi_{e}:\Sigma_{u}\to K_{b(e)}$ may be viewed as a map to the space $K_{S}$ for this $S\in\Pi$ which contains the vertex $b(e)$. Thus, for $u$ and $e$ as above, we put $\Sigma_{u}^{\Pi}:=\Sigma_{u}$ and $\varphi_{e}^{\Pi}:=\varphi_{e}$. Now, viewing any subtree $S\in\Pi$ as a vertex of $T^{\Pi}$, for any vertex $u\in N^{T^{\Pi}}_{S}$ denote by $t^{S}_{u}$ this vertex of the subtree $S$ which is adjacent to $u$ in $T$, and note that, as a consequence of Fact 2.6, the family
@@
The following result has been proved as Theorem 3.A.1 in [23].
@@
For any tree system $\Theta$ of metric compacta, and its any consolidation $\Theta^{\Pi}$, the limits $\lim\Theta$ and $\lim\Theta^{\Pi}$ are canonically homeomorphic.
@@ 3. The reduced Bowditch JSJ splitting and rigid cluster factors
In this section we introduce a new concept of the reduced Bowditch JSJ splitting of a 1-ended hyperbolic group, which is a variation of the classical Bowditch JSJ splitting. In particular, we describe the factors of this splitting, especially the so called rigid cluster factors which appear in the statement of our main result—Theorem A of the introduction. We start with recalling basic features of the classical Bowditch JSJ splitting.
@@ 3.1. The Bowditch JSJ splitting of a 1-ended hyperbolic group
We refer the reader to [1] or to Section 5 in [15] for a detailed exposition of the Bowditch JSJ splitting. Here we only mention briefly that this splitting describes in some clever way the configuration of all local cutpoints in the Gromov boundary $\partial G$ of a 1-ended hyperbolic group $G$, and relates it to certain canonical JSJ splitting of $G$ (so that the Bowditch JSJ tree appearing as part of the data of this decomposition is canonically identified with the Bass-Serre tree of the corresponding JSJ splitting of $G$).
More precisely, let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian (i.e. its Gromov boundary $\partial G$ is not homeomorphic to the circle $S^{1}$). Recall that, under these assumptions, $\partial G$ is a connected and locally connected compact metrizable space without global cutpoints [1]. Moreover, each local cutpoint $p$ of $\partial G$ has finite degree (where degree of $p$ is, by definition, the number of ends of a locally compact space $\partial G\setminus\{p\}$). The set of all local cutpoints of $\partial G$ is split into classes, so that any two points from the same class form a cut pair, any two points in the same class have the same degree and that this degree is equal to the number of connected components in the complement of the corresponding cut pair. This splitting into classes is natural, in the sense that classes are mapped to classes by homeomorphisms of $\partial G$. Each class is either infinite (in which case it consists of points of degree 2) or it is of size 2 (i.e. it consists of precisely 2 points). Two classes $\zeta_{1}$ and $\zeta_{2}$ are separated if they lie in distinct components of the complement $\partial G\setminus\{x,y\}$ of a cut pair $x,y\in\eta$ from some other class $\eta$. An object called the Bowditch JSJ tree for $G$, which we denote by ${T}_{G}$, has vertices $v$ (accompanied with some quasi-convex subgroups $G(v)<G$) of the following four types:
@@
    vertices $v$ related to infinite classes of local cutpoints of degree 2, the so called necklaces, for which the associated groups $G(v)$ are virtually free and coincide with the stabilizers of the corresponding necklaces under the action of $G$ on $\partial G$; the boundary $\partial G(v)\subset\partial G$ coincides then with the closure of the corresponding necklace, and it is homeomorphic to the Cantor set; such vertices $v$ have countable infinite order in ${T}_{G}$;
@@
    vertices $v$ related to classes of size 2, for which the associated groups $G(v)$ are equal to the $G$-stabilizers of the corresponding classes; these groups $G(v)$ are maximal 2-ended subgroups of $G$ and their boundaries $\partial G(v)$ coincide with the corresponding classes; the order of such a vertex $v$ in the tree ${T}_{G}$ is equal to the order of any local cutpoint in the corresponding class, and it is an integer $\geq 2$;
@@
    vertices $v$ related to infinite maximal families of pairwise not separated classes called stars, for which the associated groups $G(v)$ are equal to the $G$-stabilizers of unions of the corresponding stars; the order of such a vertex $v$ in the tree ${T}_{G}$ is equal to the cardinality of the family of classes in the corresponding star, and it is infinite countable;
@@
    vertices $v$ corresponding to pairs consisting of a necklace $C$ and a star $S$ containing this necklace; if we denote by $v_{1}$ and $v_{3}$ the vertices corresponding to the necklace $C$ and to the star $S$, respectively, then $G(v)=G({v_{1}})\cap G({v_{3}})$, and it is a maximal 2-ended subgroup of $G$; the order of such a vertex $v$ in the tree ${T}_{G}$ is equal to 2, and the boundary $\partial G(v)$ is contained in $C$.
@@
    an edge $e$ connecting a pair of vertices $v_{1},v_{2}$ of types (v1) and (v2), respectively; such an edge appears if and only if the class of size 2 corresponding to $v_{2}$ is contained in the closure of the infinite class corresponding to $v_{1}$; moreover, in this case the class corresponding to $v_{2}$ consists of local cutpoints of orders $\geq 3$;
@@
Let $v$ be a black vertex of ${T}_{G}$, i.e. either a rigid or a flexible vertex. Put $K_{v}:=\partial G(v)$, and for any vertex $u$ of ${T}_{G}$ adjacent to $v$, $u\in N_{v}$, put $\Sigma_{u}:=\partial G(u)=\partial G([u,v])$. Then the pair $(K_{v},\{\Sigma_{u}:u\in N_{v}\})$ is a space with peripherals. Moreover, if $v$ is flexible then this space with peripherals is (homeomorphic to) the punctured circle $(S^{1})^{\circ}$.
We now turn to the description of certain tree system canonically associated to the Bowditch JSJ splitting of $G$, and to the expression of the Gromov boundary of $G$ as the limit of this tree system. Denote by $V_{{T}_{G}}^{b}$ and $V_{{T}_{G}}^{w}$ the sets of all black and white vertices of the tree ${T}_{G}$, respectively. Consider the following tree system $\Theta=\Theta(G)$:
@@
[…]
@@
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian, and let $\Theta(G)$ be the associated tree system. Then the Gromov boundary $\partial G$ is naturally homeomorphic to the limit of the tree system $\Theta(G)$, i.e.

### v2

@@ 2.3. Limit of a tree system
We now describe, following the exposition in Section 1 of [Ś20], an object called the limit of a tree system $\Theta=(T,\{K_{t}\},\{\Sigma_{u}\},\{\varphi_{e}\})$, denoted $\lim\Theta$, starting with its description as a set. Denote by #$\#\Theta$ the quotient
@@
Again, for any finite b-subtree $F$ of $T$ let $R_{F}$ be the set of all rays $\rho$ in $T$ with $\rho(0)\in N_{F}$, and with $\rho$ disjoint from $F$. Note that the map $\rho\to[\rho]$ is then a bijection from $R_{F}$ to the set $\partial T$ of all ends of $T$. Given a subset $U\subset K_{F}$ as above (i.e. saturated with respect to $\Omega_{F}$), put
@@
where all summands in the above union are viewed as subsets of $\lim\Theta$; clearly, ${\mathcal{G}}(U)$ is then also a subset in $\lim\Theta$. We consider the topology in the set $\lim\Theta$ given by the basis $\mathcal{B}$ consisting of all sets ${\mathcal{G}}(U)$, for all finite b-subtrees $F$ in $T$, and all open subsets $U\subset K_{F}$ saturated with respect to $\Omega_{F}$. We leave it as an easy and instructive exercise to check that the family $\mathcal{B}$ is closed under finite intersections, and hence it satisfies the axioms for a basis of a topology. Moreover, we have the following result, which is a slight extension of Proposition 1.C.1 of [Ś20], and which can be justified by the same proof as the latter proposition.
@@ 2.4. Functoriality with respect to subtrees
Next observation follows by a direct argument (compare the proof of Claim 1 inside the proof of Theorem 3.A.1 in [Ś20]).
@@
In the case when $S=F$ is a finite b-subtree of $T$, it is not hard to see that the limit $\lim\Theta_{F}$ coincides with the partial union $K_{F}$. It follows then from Fact 2.5 that each of the partial unions $K_{F}$ naturally embeds (as a topological space) in $\lim\Theta$. In particular, each of the constituent spaces $K_{t}$ naturally embeds in $\lim\Theta$. A similar observation shows that each of the peripheral spaces $\Sigma_{u}$ naturally embeds in $\lim\Theta$. The following result is proved as Proposition 1.C.1 in [Ś20].
@@ 2.5. An isomorphism of tree systems
Let $\Theta=(T,\{K_{t}\},\{\Sigma_{u}\},\{\varphi_{e}\})$ and $\Theta^{\prime}=(T^{\prime},\{K^{\prime}_{t}\},\{\Sigma^{\prime}_{u}\},\{\varphi^{\prime}_{e}\})$ be two tree systems of metric compacta. An isomorphism $F:\Theta\to\Theta^{\prime}$ is a tuple
$F=(\lambda,\{f_{t}:t\in V^{b}_{T}\},\{g_{u}:u\in V^{w}_{T}\})$
such that:
@@
If $\Theta,\Theta^{\prime}$ are isomorphic tree systems of metric compacta then their limits $\lim\Theta$ and $\lim\Theta^{\prime}$ are homeomorphic. More precisely, an isomorphism $F:\Theta\to\Theta^{\prime}$ induces (in the obvious way) a homeomorphism $\lim\Theta\to\lim\Theta^{\prime}$.
@@ 2.6. An inverse system associated to a tree system
be a tree system. For any finite b-subtree $F\subset T$, consider the family $\Omega_{F}$ of peripheral subspaces of the partial union $K_{F}$, as described in Section 2.3, and denote by $K_{F}^{*}=K_{F}/\Omega_{F}$ the quotient of $K_{F}$ in which all the sets from $\Omega_{F}$ are shrunk to points. More precisely, let ${\mathcal{D}}_{F}$ be the decomposition of $K_{F}$ consisting of the sets of $\Omega_{F}$ and the singletons from the complement $K_{F}\setminus\bigcup\Omega_{F}$, and let $K_{F}^{*}=K_{F}/{\mathcal{D}}_{F}$. Since the family $\Omega_{F}$ is null, the decomposition ${\mathcal{D}}_{F}$ is upper semicontinuous, and hence the quotient $K_{F}^{*}$ is metrizable (see [Dav86], Proposition 3 on page 14 and Proposition 2 on page 13). Consequently, $K_{F}^{*}$ is a metric compactum, and we call it the reduced partial union of $\Theta$ related to $F$. We denote by $q_{F}:K_{F}\to K_{F}^{*}$ the quotient map resulting from the above description of $K_{F}^{*}$.
@@
Next result is a slight extension of Proposition 1.D.1 in [Ś20], and its proof is the same as the proof of that proposition (hence we omit it).
@@ 2.7. Consolidation of a tree system
Let $\Theta=(T,\{K_{t}\},\{\Sigma_{u}\},\{\varphi_{e}\})$ be a tree system of metric compacta. Let $\Pi$ be a partition of a tree $T$ into b-subtrees, i.e., a family of b-subtrees $S\subset T$ such that the black vertex sets $V_{S}^{b}:S\in{\Pi}$ are pairwise disjoint and cover all of $V_{T}^{b}$. We allow that some of the subtrees $S\in{\Pi}$ are just single black vertices of $T$.
@@
described as follows. As a set $V^{b}_{T^{\Pi}}$ of black vertices of $T^{{\Pi}}$ we take the family $\Pi$, and as a set $V^{w}_{T^{\Pi}}$ of white vertices we take the set of those white vertices $u\in V^{w}_{T}$ which remain outside all of the subtrees $S\in\Pi$. As the edge set $E_{T^{\Pi}}$ we take the set $\{e\in E_{T}:e\notin\cup_{S\in{\Pi}}\,E_{S}\}$, so that each edge $e$ from this set is meant to connect the white vertex $w(e)\in V^{w}_{T^{\Pi}}$ with the subtree $S\in\Pi$ (viewed as a black vertex of $T^{\Pi}$) which contains the vertex $b(e)$ of $T$. Intuitively speaking, the above described tree $T^{\Pi}$ is a “dual tree” of the partition $\Pi$.
To describe the constituent spaces $K_{S}$ of the tree system $\Theta^{\Pi}$, for any subtree $S\in{\Pi}$ denote by $\Theta_{S}$ the restriction of $\Theta$ to $S$, and put $K_{S}:=\lim\Theta_{S}$. To describe the peripheral spaces $\Sigma_{u}^{\Pi}$ and the connecting maps $\varphi_{e}^{\Pi}$, note that for any vertex $u\in V^{w}_{T^{\Pi}}$ (viewed as a white vertex of $T$) and any edge $e$ with $w(e)=u$, the map $\varphi_{e}:\Sigma_{u}\to K_{b(e)}$ may be viewed as a map to the space $K_{S}$ for this $S\in\Pi$ which contains the vertex $b(e)$. Thus, for $u$ and $e$ as above, we put $\Sigma_{u}^{\Pi}:=\Sigma_{u}$ and $\varphi_{e}^{\Pi}:=\varphi_{e}$. Now, viewing any subtree $S\in\Pi$ as a vertex of $T^{\Pi}$, for any vertex $u\in N^{T^{\Pi}}_{S}$ denote by $t^{S}_{u}$ this vertex of the subtree $S$ which is adjacent to $u$ in $T$, and note that, as a consequence of Fact 2.6, the family
@@
The following result has been proved as Theorem 3.A.1 in [Ś20].
@@
For any tree system $\Theta$ of metric compacta, and any consolidation $\Theta^{\Pi}$ of $\Theta$, the limits $\lim\Theta$ and $\lim\Theta^{\Pi}$ are canonically homeomorphic.
@@ 3. The reduced Bowditch JSJ splitting and rigid cluster factors
In this section we introduce a new concept of the reduced Bowditch JSJ splitting of a 1-ended hyperbolic group, which is a variation of the classical Bowditch JSJ splitting. In particular, we describe the factors of this splitting, especially the so called rigid cluster factors which appear in the statement of our main result, Theorem A of the introduction. We start with recalling basic features of the classical Bowditch JSJ splitting.
@@ 3.1. The Bowditch JSJ splitting of a 1-ended hyperbolic group
We refer the reader to [Bow98] or to Section 5 in [KW22] for a detailed exposition of the Bowditch JSJ splitting. Here we only mention briefly that this splitting describes in some clever way the configuration of all local cutpoints in the Gromov boundary $\partial G$ of a 1-ended hyperbolic group $G$, and relates it to a certain canonical JSJ splitting of $G$ (so that the Bowditch JSJ tree appearing as part of the data of this construction is canonically identified with the Bass-Serre tree of the corresponding JSJ splitting of $G$).
More precisely, let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian (i.e. its Gromov boundary $\partial G$ is connected and not homeomorphic to the circle $S^{1}$). Recall that, under these assumptions, $\partial G$ is a connected and locally connected compact metrizable space without global cutpoints [Bow98]. Moreover, each local cutpoint $p$ of $\partial G$ has finite degree (where degree of $p$ is, by definition, the number of ends of a locally compact space $\partial G\setminus\{p\}$). The set of all local cutpoints of $\partial G$ is split into classes, so that any two points from the same class form a cut pair, any two points in the same class have the same degree and that this degree is equal to the number of connected components in the complement of the corresponding cut pair. This splitting into classes is natural, in the sense that classes are mapped to classes by homeomorphisms of $\partial G$. Each class is either infinite (in which case it consists of points of degree 2) or it is of size 2 (i.e. it consists of precisely 2 points). Two classes $\zeta_{1}$ and $\zeta_{2}$ are separated if they lie in distinct components of the complement $\partial G\setminus\{x,y\}$ of a cut pair $x,y\in\eta$ from some other class $\eta$. An object called the Bowditch JSJ tree for $G$, which we denote by ${T}_{G}$, has vertices $v$ (accompanied with some quasi-convex subgroups $G(v)<G$) of the following four types:
@@
    vertices $v$ related to infinite classes of local cutpoints of degree 2, the so called necklaces, for which the associated groups $G(v)$ are virtually free and coincide with the stabilizers of the corresponding necklaces under the action of $G$ on $\partial G$; the boundary $\partial G(v)\subset\partial G$ coincides then with the closure of the corresponding necklace, and it is homeomorphic to the Cantor set; such vertices $v$ have countable infinite degree in ${T}_{G}$;
@@
    vertices $v$ related to classes of size 2, for which the associated groups $G(v)$ are equal to the $G$-stabilizers of the corresponding classes; these groups $G(v)$ are maximal 2-ended subgroups of $G$ and their boundaries $\partial G(v)$ coincide with the corresponding classes; the degree of such a vertex $v$ in the tree ${T}_{G}$ is equal to the degree of any local cutpoint in the corresponding class, which is an integer $\geq 2$;
@@
    vertices $v$ related to infinite maximal families of pairwise not separated classes called stars, for which the associated groups $G(v)$ are equal to the $G$-stabilizers of unions of the corresponding stars; the degree of such a vertex $v$ in the tree ${T}_{G}$ is equal to the cardinality of the family of classes in the corresponding star, which is always infinite countable;
@@
    vertices $v$ corresponding to pairs consisting of a necklace $C$ and a star $S$ containing this necklace; if we denote by $v_{1}$ and $v_{3}$ the vertices corresponding to the necklace $C$ and to the star $S$, respectively, then $G(v)=G({v_{1}})\cap G({v_{3}})$, which is a maximal 2-ended subgroup of $G$; the degree of such a vertex $v$ in the tree ${T}_{G}$ is equal to 2, and the boundary $\partial G(v)$ is contained in $C$.
@@
    an edge $e$ connecting a pair of vertices $v_{1},v_{2}$ of types (v1) and (v2), respectively; such an edge appears if and only if the class of size 2 corresponding to $v_{2}$ is contained in the closure of the infinite class corresponding to $v_{1}$; moreover, in this case the class corresponding to $v_{2}$ consists of local cutpoints of degrees $\geq 3$;
@@
Let $v$ be a black vertex of ${T}_{G}$, i.e., either a rigid or a flexible vertex. Put $K_{v}:=\partial G(v)$, and for any vertex $u$ of ${T}_{G}$ adjacent to $v$, $u\in N_{v}$, put $\Sigma_{u}:=\partial G(u)=\partial G([u,v])$. Then the pair $(K_{v},\{\Sigma_{u}:u\in N_{v}\})$ is a space with peripherals. Moreover, if $v$ is flexible then this space with peripherals is (homeomorphic to) the punctured circle $(S^{1})^{\circ}$.
We now turn to the description of a certain tree system canonically associated to the Bowditch JSJ splitting of $G$, and to the expression of the Gromov boundary of $G$ as the limit of this tree system. Denote by $V_{{T}_{G}}^{b}$ and $V_{{T}_{G}}^{w}$ the sets of all black and white vertices of the tree ${T}_{G}$, respectively. Consider the following tree system $\Theta=\Theta(G)$:
@@
where each $\varphi_{e}:\partial G({w(e)})\to\partial G({b(e)})$ is the natural inclusion. We will call $\Theta(G)$ the tree system associated to the Bowditch JSJ splitting of $G$, or concisely the tree system associated to $G$. The next result follows fairly directly by the arguments provided by Alexandre Martin (see Corollary 9.19 in [Mar14] or Theorem 2.25 in [CM17]). In the proof provided below we explain briefly how to apply these results in our situation, by referring to the exposition in [CM17].
@@
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian, and let $\Theta(G)$ be the associated tree system. Then the Gromov boundary $\partial G$ is homeomorphic to the limit of the tree system $\Theta(G)$, i.e.,